OpenStudy (anonymous):

The force, F, between two particles is inversely proportional to the square of the distance between them. The force is 36 units when the distance between the particles is r metres. Find the force when the distance is 3r metres.

OpenStudy (anonymous):

You might want to know this: \(Fr^2 = k\) where k is a constant. Find k, find it all. Do you need further help?

OpenStudy (anonymous):

so 36r^2=k and f = 36/3 = 12 ? but the answer frm my assessment bk is 4 .. please help again, thank you ^^

OpenStudy (anonymous):

Well, You might notice that it's \(r^2\) so it's \(f 9r^2 = k\)

OpenStudy (anonymous):

oh so it is 36/9 = 4 instead ! Thank you so much ^^

OpenStudy (anonymous):

You're welcome :)